OpenStudy (anonymous):

someone take a look at this please lim_{x rightarrow -1} (x^2 - 7x + 8/ x - 5)^2

OpenStudy (anonymous):

\[\lim_{x \rightarrow -1} (x^2 - 7x + 8/ x - 5)^2\]

myininaya (myininaya):

plug in -1 since the function is continuous at -1

OpenStudy (anonymous):

i tried that and i dont know what im doing wrong but my software wont accept the result

myininaya (myininaya):

\[\frac{(-1)^2-7(-1)+8}{(-1-5)^2}\] is that right?

myininaya (myininaya):

or is the square on the whole thing?

OpenStudy (anonymous):

what happens to the exponent outside the paranthesis

myininaya (myininaya):

could you answer my question first? is the square on the whole thing or just the bottom?

OpenStudy (anonymous):

the whole thing

myininaya (myininaya):

\[(\frac{(-1)^2-7(-1)+8}{(-1-5)})^2\]

myininaya (myininaya):

so perform what is inside first

myininaya (myininaya):

\[(\frac{1+7+8}{-6})^2=(\frac{16}{-6})^2=(\frac{-8}{3})^2=\frac{64}{9}\]

OpenStudy (anonymous):

i see. thank you very much. i have alot of similar problems like this. would you be willing to work with me please?

myininaya (myininaya):

just post the questions as a new thread and i will work a little longer tonight but not much more

OpenStudy (anonymous):

ok thanks